A 200 N force pulls a 100 N box on a frictionless table. The acceleration of the block is ___ m/s/s

If instead, a 200 N weight was hung on the string then the acceleration would be ___ m/s/s
Answer #1

Net force acting on the block when it is on the table = 200N

acceleartion of the block = F/m = F/(W/g)

where W is the weight of the block

a = 200/(100/9.8)

a = 19.6 m/s2

when the block is hung from the string

net force = weight + F(applied)

F(net) = 100 + 200 = 300N

acceleration when the block is hung with a string = 300/(100/9.8)

            = 29.4 m/s2